OpenStudy (anonymous):

Use the properties of exponents to simplify each of the following expressions, writing your answer with only positive exponents.

OpenStudy (anonymous):

\[\left[ 5m^4n^{-2})^{-1} \right]^{-2}\]

OpenStudy (anonymous):

(m^8)/(25n^4)

OpenStudy (anonymous):

How did you solve it Meverett04? Can you show me the steps please?

OpenStudy (escherichiarinku):

\[(x^m)^n = x^{mn}\] and \[x^{-m} = \frac{1}{x^m}\] is what you need here

OpenStudy (anonymous):

For m, multiply the exponents 4 times -1 times -2 equals 8

OpenStudy (anonymous):

For n, multiply the exponents -2 times -1 times -2 equals -4 but since we have a negative four, this will go in the denominator with a positive 4

OpenStudy (anonymous):

For 5 5^2 or five squared is 25, but since we have a negative exponent, we place the 25 in the denominator

OpenStudy (anonymous):

thanks man!

OpenStudy (anonymous):

here is another problem if you would like to help me solve it.

OpenStudy (anonymous):

\[\left[ 100^0(x^{-1}y^3)^{-1}/x^2y \right]^{-3}\]

OpenStudy (anonymous):

I didn't know how to post it as a fraction but if i did x^2y would be the denominator.